OpenStudy (sleepyjess):

Find the derivative of f(x) = \(-\dfrac 7x\) at x = -3.

OpenStudy (sleepyjess):

@mathslover

OpenStudy (sleepyjess):

Would I use the power rule?

mathslover (mathslover):

Yup, 7 is a constant. So, it has nothing to do with differentiation. Just keep it out for once, and differentiate 1/x

mathslover (mathslover):

Oh and yes, don't forget that "-1" . \(\cfrac{d}{dx} \left(\cfrac{-7}{x}\right) = -7 \cfrac{d}{dx} \left( \cfrac{1}{x}\right) = -7 \cfrac{d}{dx} \left(x^{-1}\right) \) This is what I mean.

mathslover (mathslover):

Just use the power rule now : \(\cfrac{d}{dx} x^n = nx^{n-1} \) Let me know what you get.

OpenStudy (sleepyjess):

\(-7x^{-2}\)?

mathslover (mathslover):

Check the sign! (in front of 7)

OpenStudy (sleepyjess):

7x^-2 ?

mathslover (mathslover):

Yep! Now plug-in x = -3

OpenStudy (sleepyjess):

wait, isn't there a rule or something for negative exponents?

mathslover (mathslover):

No, there isn't any specific, as far as my knowledge is concerned. http://www.wolframalpha.com/input/?i=differentiate+-7%2Fx The answer is same, so, don't worry, you're going right.

OpenStudy (sleepyjess):

7(-3)\(^{-2}\)

mathslover (mathslover):

Yeah, that's correct! Good work.

OpenStudy (sleepyjess):

:D what's after that?

rvc (rvc):

find it :)

mathslover (mathslover):

Uhmm.. simplify? And then party? :O :P

OpenStudy (sleepyjess):

xD

OpenStudy (sleepyjess):

7/9!

rvc (rvc):

jess i have a tutorial on derivatives you can just revise that ;)

OpenStudy (sleepyjess):

:O

OpenStudy (sleepyjess):

I didn't know that!

rvc (rvc):

correct All The Best :)

OpenStudy (sleepyjess):

Is the post showing closed for you guys?

rvc (rvc):

yes